OpenStudy (camerondoherty):

\[f(x) = -16x^2 + 60x + 16\] Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work.

OpenStudy (jtvatsim):

First part of the answer is clear. Second, can be found quickly with Calculus.

OpenStudy (camerondoherty):

Calculus? I'm only in Alg 1. Haha

OpenStudy (jtvatsim):

But Calculus makes the world go round. :) It's still possible to solve. Let's see what the others suggest.

OpenStudy (camerondoherty):

And I have no idea of the first part It may be clear to you but its definately not clear to me cx

OpenStudy (jtvatsim):

haha... I know, I was assuming too much. :P

OpenStudy (jtvatsim):

Do you know what shape the graph will be?

OpenStudy (triciaal):

standard form ax^2 + bx + c = 0 vertex form a(x-h)^2 + k = 0 x-coordinate of the vertex h = -b/2a y value for vertex = result for f(h)

OpenStudy (camerondoherty):

Wait, so... what is a, b, c, x, and k? cx

OpenStudy (camerondoherty):

Like, do I sub in anything

OpenStudy (triciaal):

since you are not ready for calculus, pick 2 points one on each side of the vertex and use this as the x value in f(x) and calculate the y values you will be able to do a sketch (x, y) and determine if the vertex was a maximum or minimum

OpenStudy (camerondoherty):

So, do I graph it on desmos and pick two points?

OpenStudy (triciaal):

do you know how to factorize the equation given?

OpenStudy (camerondoherty):

Yep

OpenStudy (camerondoherty):

I actually already did 4(x - 4)(4x + 1)

OpenStudy (camerondoherty):

before

OpenStudy (triciaal):

the vertex of the equation is when the function = 0 same as saying when y = 0 find the x-values

OpenStudy (camerondoherty):

I had already found those as well cx I got 4 and -1/4

OpenStudy (triciaal):

great so when y = 0 what are the x values?

OpenStudy (camerondoherty):

the zeroes

OpenStudy (triciaal):

so the points are (x, y) and y = 0

OpenStudy (camerondoherty):

(4,0) & (-1/4,0)

OpenStudy (camerondoherty):

Ohhhh, and those are the vertex coordinates?

OpenStudy (camerondoherty):

I still don't get how it is a maximum though :/

OpenStudy (triciaal):

correction to this post thanks@jtvatsim the vertex of the equation is when the function = 0 same as saying when y = 0 find the x-values the original is correct -b/2a =h the x-coordinate

OpenStudy (camerondoherty):

Wait, what? cx

OpenStudy (jtvatsim):

I'll draw the picture of where we are at right now...